PROJECT CASH FLOW Colsen Communications is trying to estimate the first-year cash flow (at Year 1) for a proposed project. The financial staff has collected the following information on the project: Sales revenues $5 million Operating costs (excluding depreciation) 3.5 million Depreciation 1 million Interest expense 1 million The company has a 40% tax rate, and its WACC is 13%. Write out your answers completely. For example, 13 million should be entered as 13,000,000. What is the project's cash flow for the first year (t = 1)? Round your answer to the nearest dollar. $ If this project would cannibalize other projects by $0.5 million of cash flow before taxes per year, how would this change your answer to part a? Round your answer to the nearest dollar. The firm's project's cash flow would now be $ . Ignore part b. If the tax rate dropped to 35%, how would that change your answer to part a? Round your answer to the nearest dollar. The firm's project's cash flow would by $ =? .
Answer #1

What is the project's cash flow for the first year (t = 1)?
=(5000000-3500000-100000)*(1-40%)+1000000
=1840000.00

If this project would cannibalize other projects by $0.5 million of cash flow before taxes per year, how would this change your answer to part a?
=(5000000-3500000-100000-500000)*(1-40%)+1000000
=1540000.00

If the tax rate dropped to 35%, how would that change your answer to part a?
=(5000000-3500000-100000)*(1-35%)+1000000
=1910000.00

Increase by 70000